# GLOBALIZATION? TRADE WAR? A COUNTERBALANCE PERSPECTIVE

### v4

@@ 2.2 Mixed Cooperation and Noncooperation
The $1\times n$ row vector $\pi$, called authority distribution in Hu and Shapley, (2003), also solves the following counterbalance equilibrium
@@
subject to the normalization conditions $\sum\limits_{i=1}^{n}\pi_{i}=1$ and $\pi_{i}\geq 0$. There exist many counterbalanced systems in the physical and human worlds, such as ecological systems, the USA’s government system, and China’s five-element theory.
The counterbalance in Eq. (3) implies $\pi$’s mixed cooperative and noncooperative properties — even the most competitive economy needs both cooperation and competition. Between China and the USA from 1980 to 2020, there was more cooperation in the first thirty years than in the last ten years. Though unfamiliar to empirical macroeconomics, this mixture could have an edge in the future (e.g., Allen, 2000, page 147). An econometric analysis, for example, could hardly capture the complexity of antagonism, where any estimated coefficient or implicated effect cannot be both significantly positive and significantly negative.
From the noncooperative aspect, as $\sum\limits_{i=1}^{n}\pi_{i}=1$, an increase of $\pi_{j}$ may imply a decrease of $\pi_{i}$. So, in theory, there are $n(n-1)/2$ potential trade conflicts, either small or large, in the trade system. Most of them are minor trade disagreements or disputes, so the “war” metaphor is hyperbolic in these cases. However, it is worth mitigating even minor disputes before they become problematic or catastrophic. Also, the increase of $\pi_{j}$ is not evenly redeemed by other economies; they often wax and wane with some common causality initialized by changes in the matrix $P$. In Sections 3 and 5, we distinguish the first move off the diagonal and that on the diagonal of $P$, respectively.
From the cooperative aspect, country $i$ should help $j$ improve $\pi_{j}$ whenever $P_{ji}>0$ because $\pi_{j}P_{ji}$ is part of $\pi_{i}$, as shown in
@@
from Eq. (3). Also, Eq. (3) implies that $\pi=\pi P=\pi P^{2}=\cdots=\pi P^{t}$ in which
@@
for any $t=1,2,3,\cdots.$ When $\pi_{i}>0$, as $\lim\limits_{t\to\infty}\left(P^{t}\right)_{ji}=\pi_{i}>0$, therefore, $\left(P^{t}\right)_{ji}>0$ for all large $t$ and all $j\in\mathcal{N}$. So, Eq. (5) further implies that $i$ should help all other countries in the trade system, including the poorest and the least competitive ones, to enhance their $\pi_{j}$. Thus, the double-edged counterbalance implies that a carrot-and-stick approach would better leverage country $i$’s position in the trade system when dealing with its trade partner $j$. The trade-off is how much it should assist or contest $j$ without sacrificing its national interests and even increasing $\pi_{i}$.
@@
We can physically visualize $\pi_{i}$ as a container or pool in the dynamics of counterbalance equilibrium, in which comparative advantages are the primary driving force for the flows. On the one hand, in the inflow equation Eq. (4), $i$ accumulates or absorbs authority from other countries through its direct influences. It derives more authority from influential trade partners (i.e., $\pi_{j}$ is large) than from non-influential ones, other things being equal. It also derives more authority from countries on which it has a sizable direct influence (i.e., $P_{ji}$ is large), other things being equal. In Figure I, country $i$ imports goods from $j$ through the $i$th column of $P$ and so, authority flows from $j$ to $i$ alongside the same column. On the other hand, $i$ also contributes to other countries as seen in the following outflow equation
$\pi_{j}=\pi_{i}P_{ij}+\sum\limits_{k\not=i}\pi_{k}P_{kj}.$
The larger $P_{ij}$ or $\pi_{i}$, the more $i$ contributes to $j$’s authority $\pi_{j}$, other things being equal. In Figure I, country $i$ redistributes $\pi_{i}$ through the $i$th row of $P$, according to its shares of exports.
In the accounting of $\pi$, the two countries have a break-even point when
@@
Intuitively, if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, then country $i$ takes more from or gives more to country $j$, respectively. For country $i$, a simple implication of Eq. (6) is to decrease or increase its imports from $j$ if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, respectively. It can also adjust its exports to $j$ accordingly. However, the outflow from $i$ to $j$ does not necessarily match the inflow. With involvements of third parties, the inflows and outflows meet at the equilibrium Eq. (3) for $i$, i.e. $\pi_{i}P_{ij}+\sum\limits_{k\not=j}\pi_{i}P_{ik}=\pi_{j}P_{ji}+\sum\limits_{k\not=j}\pi_{k}P_{ki}=\pi_{i}$. For the same reason, it could be overly restrictive for all countries to balance their bilateral trades. For example, country $i$ still maintains an overall zero net balance when it has a trade surplus with $j$ and a trade deficit with $k$ of the same amount.
2.3 $\pi_{i}$ as Competitive Advantage
@@
Country $i$’s competitive advantage refers to its ability to do better than its competitors in the trade system. For any product exported from country $i$ to $j$, $i$’s competitors include all countries other than itself. So, its competitive advantage extends from bilateral country relations to a relationship with the whole system — it is the ability to beat the best of all other countries in the $j$th market for any $j\in\mathcal{N}$. On the contrary, comparative advantage is an economy’s leverage of specialization, resource endowment, and technology over a specific trade partner on particular services or products; it is a bilateral relationship. According to the literature (e.g., Porter, 1985; Stutz and Warf, 2010), competitive advantages rest on the direct control of competitive factors of production, such as access to rare resources or advanced technology, cheap but highly skilled labor, and a low cost of capital. These factors, imported from the outside or manufactured domestically, are assembled in the final products. Thus, the imported goods carry the original producer’s competitiveness and should be attributed to the original maker — in Eq. (4), $\pi_{j}P_{ji}$ contains $j$’s return of competitiveness to $\pi_{i}$. In the above smartphone example, the competitive factors include the chips and patents from the USA, the cheap labor from China, the precision manufacturing in Japan, and the sophisticated cameras from Korea. Various studies have also described qualitative strategies to enhance firm competitiveness. For example, Porter (1985) lists three generic strategies: cost leadership, differentiation, and focus.
@@
We find that $\pi$ quantifies national competitiveness. To introduce the quantitative competitiveness, we ignore the home bias in trade (cf., McCallum, 1995) and make a strong assumption for the time being: no bilateral trade deficit nor surplus. Denote $\tau_{j}$ the $j$th row of $P$. Because of no bilateral surplus nor deficit, $\tau_{j}$ except for its $j$th element equals the imports from other countries to $j$, all in fractions of $j$’s production. So, it measures other countries’ aggregate comparative advantages over country $j$. Besides, all other countries also compete for the comparative advantages over $j$; a large value in $\tau_{j}$ means a solid competitiveness to occupy the $j$th market and win a lion share of $j$’s imports. If we use the $j$th market as the reference to rank the competitiveness of all countries, then the competitiveness score is the vector $\tau_{j}$, which has a unit sum. Yet, the $j$th value is not excluded from $\tau_{j}$ because domestic products also compete with country $j$’s imports when the consumers make their consumption choices, and they have no home bias. We apply endogenous weighting to the references to combine these scores $\tau_{1},\tau_{2},\cdots,\tau_{n}$. The weights are not exogenously assigned but endogenously determined. Also, more weight should be placed on a more competitive reference market than on a less competitive one because it is more challenging to occupy a competitive market than a less competitive one. Let the unknown weights be $\rho=(\rho_{1},\rho_{2},\cdots,\rho_{n})$ with $\sum\limits_{j=1}^{n}\rho_{j}=1$ and $\rho_{j}\geq 0$ for all $j\in\mathcal{N}$. Then, we have a weighted competitiveness score $\sum\limits_{j=1}^{n}\rho_{j}\tau_{j}=\rho P$. So, both the weighted score and the weight vector quantify the competitiveness of the countries. To be consistent, the weighted average $\rho P$ should be a positive multiple of the weight vector $\rho$. By Theorem 1, $\rho$ is just $\pi$.
@@
The national competitiveness $\pi_{i}$ is not the sum of the individual competitiveness of the country’s domestic firms, though they are positively associated. In the endogenous weighting, worldwide consumers unconsciously participate in the evaluation of competitiveness, and all products are involved. Essentially, countries unknowingly judge countries themselves without external interventions from business interests, mainstream media, advertising, governments, or even military threats. The structure of the global trade network is, to a great extent, endogenously determined by the distribution of firm productivity in individual countries. Nevertheless, government policies can affect the distribution without significantly mitigating the welfare of international trade by switching imports or exports from one country to another. It can also enhance firm productivity through tax exemption, subsidies, and heavy tariffs on foreign products.
@@
Also, $\pi$ is a von Neumann-Morgenstern Bergson social welfare function (e.g., Harsanyi, 1955; Hammond, 1987) since it is a positively weighted sum of von Neumann-Morgenstern individual utility functions $\tau_{1},\tau_{2},\cdots,\tau_{n}$; the weights are $\pi$ itself. Thus, the weighted utility function $\pi$ is not determined by how much production factors are consumed but by how much the consumers prefer the products, judged regardless of national interests or preferences by individual countries. Pareto optimality is satisfied when weighing these individual utilities: $\pi_{j}\geq\pi_{k}$ if $P_{ij}\geq P_{ik}$ for all $i\in\mathcal{N}$. The optimality remains invariant if any individual utility function is scaled by a positive constant or added by a constant.
@@
However, the welfare function $\pi$ departs significantly from widely-used utility or welfare functions in the literature on trade wars or globalization (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014). It is not explicitly expressed in a specific functional form of production factors but implied from the counterbalance equilibrium because indirect exports also act upon the final products or services. Any exported goods and services from one country to another have different utilities for the exporting and importing countries. Therefore, each country has its unique utility function, and a single utility function or linear regression cannot capture the diversity of all countries. Also, we use fewer assumptions to accomplish $\pi$ by ignoring how efficiently each country packages these factors into final products and how differently it trades the products with others. It is precisely derived from the data and uses no parametric models. Thus, no unknown parameters are to be estimated, and no residuals exist to differentiate the data from a parametric model. According to Barney and Felin, (2013), micro-foundations research is still heavily debated, with management, strategy, and organization scholars having varying views on the link between macroeconomic phenomena and microeconomic agents’ behaviors. Because of broad firm and national heterogeneity, micro-founded strategic movements are particularly challenging on the network with flows of millions of products over tens of thousands of directional edges. With this complexity, a certain level of abstraction and ambiguity is necessary to grasp the essential parts of the network. Also, unjustified assumptions and overly mathematization could mislead the conclusions and policy implications (e.g., Romer, 2015).
@@
The index $\pi_{i}$ contains an inalienable part besides absorbing spillover effects through $P$. There are $t-1$ occurrences of $P_{ii}$ in Eq. (2), which largely accounts for $\pi_{i}P_{ii}$ in Eq. (4) as $t\to\infty$. A global value chain generally takes only finite steps, but it also contains indefinite domestic loops when we trace all the added values of the final products. In the smartphone example, the manufacturing labor’s education, skills, and experience are ultimately achieved domestically over many years, but all somehow contribute to the final product. Also, many USA workers develop the patents and design software domestically for an extended period. These inner loops affect other countries but not vice versa; this part of competitiveness never flows in from external. Thus, social infrastructure and political institutions, monetary and fiscal policies, microeconomic environments, and other quality determinants that help do business all contribute to national competitiveness (e.g., Delgado et al., 2012). Also, raw oil fields, natural gas reserves, rare minerals, and patented technologies do not come from external sources and remain with $\pi_{i}P_{ii}$ as the nation’s core competitive factors. Refraining exports of rare earth elements or licensed foreign usage of patented technologies, for example, could destroy one or more global value chains.
@@
For a specific global value chain, on the other end of the inalienable parts is the design and assembly of the final product. Product leadership bridges the market, the consumers, strategic direction, technology innovation, sales, and production factors together; this leadership is another type of competitiveness. Between the end product and the raw resources in the chain are a series of value-added processes for intermediate goods. Their competitiveness lies in cost efficiency or cost leadership, including the cost of financing. Technological advancements reduce labor costs, and public safety and social stability curtail the risk of costs.
@@
Finally, Eq. (2) implies a solid positive correlation between $\pi_{i}$ and $g_{i}$, the breadth of the economy. Nations sell goods that they have in abundance and buy goods that are in short supply domestically. With great abundance and shortage, country $i$ could have vast nonzero elements in the $i$th column and row of the matrix $P$, resulting in astronomical nonzero items summed in Eq. (2). A large economy also produces a high differentiation of goods and services, which bear location advantage and minimal transaction costs for domestic consumers. Free of tariffs and overseas shipping, they also contain a price advantage. For these reasons, $P_{ii}$ is often the largest in the $i$th row of $P$, with or without the home bias. Therefore, a massive economy with a comprehensively integrated industrial system (such as China and the USA) is likely, but not necessary, to have a large value in $\pi$. A country with a single industrial structure is vulnerable to brutal competition, and its $\pi_{i}$ is fragile and volatile when battling with the USA for competitiveness.
@@
Therefore, we use $\pi_{i}$ as an objective for country $i$ when making trade policies. It quantifies the economy’s depth and interplays with breadth; it also incorporates cooperation and non-cooperation with others. It is endogenously extracted from the comparative advantages of all products and services. As each country has a unique location, natural and human resources, and economic systems, a single utility function, which presumes one stereotype for all economies, may be inadequate to capture this diversity. In our study, we recognize that each country, as is its trade stance with partners, is different. Each adds inclusiveness value to the global economy, of which trade between nations is a mainstay.

### v5

@@ 2.2 Mixed Cooperation and Noncooperation
The $1\times n$ row vector $\pi$, referred to as the authority distribution in Hu and Shapley (2003), also solves the following counterbalance equilibrium:
@@
subject to the normalization conditions $\sum\limits_{i=1}^{n}\pi_{i}=1$ and non-negativity condition $\pi_{i}\geq 0$. Numerous counterbalanced systems exist in both the physical and human worlds, such as ecological systems, the USA’s government system, and China’s five-element theory.
The counterbalance in Eq. (3) implies $\pi$’s mixed cooperative and noncooperative properties. This is a distinctive feature between comparative and competitive advantages; while both promote collaborations, competitive advantage may bring trade wars and deglobalization. For instance, between China and the USA from 1980 to 2020, there was more cooperation in the first thirty years than in the last ten years. Although unfamiliar to empirical macroeconomics, this mixture could have an edge in the future (e.g., Allen, 2000, page 147). Econometric analysis often struggles to capture the complexity of antagonism, where any estimated coefficient or implicated effect cannot be both significantly positive and significantly negative.
From a non-cooperative perspective, as $\sum\limits_{i=1}^{n}\pi_{i}=1$, an increase in $\pi_{j}$ may imply a decrease in $\pi_{i}$. Thus, theoretically, there are $n(n-1)/2$ potential trade conflicts, either small or large, within the trade system. Most of these are minor trade disagreements or disputes, making the “war” metaphor hyperbolic in these cases. However, it is worth mitigating even minor disputes before they become problematic or catastrophic. Additionally, the increase in $\pi_{j}$ is not evenly redeemed by other economies; they often wax and wane with some common causality initiated by changes in the matrix $P$. In Sections 3 and 5, we distinguish the first move off the diagonal and that on the diagonal of $P$, respectively, for bilateral trade wars and globalization.
From a cooperative perspective, country $i$ should assist country $j$ in improving $\pi_{j}$ whenever $P_{ji}>0$, as $\pi_{j}P_{ji}$ is part of $\pi_{i}$, as shown in:
@@
from Eq. (3). Eq. (3) also implies that $\pi=\pi P=\pi P^{2}=\cdots=\pi P^{t},$ where
@@
for any $t=1,2,3,\cdots.$ When $\pi_{i}>0$, as $\lim\limits_{t\to\infty}\left(P^{t}\right)_{ji}=\pi_{i}>0$, therefore, $\left(P^{t}\right)_{ji}>0$ for all large $t$ and all $j\in\mathcal{N}$. Thus, Eq. (5) further implies that country $i$ should assist all other countries in the trade system, including the poorest and least competitive ones, to enhance their $\pi_{j}$. Hence, the double-edged counterbalance suggests that a carrot-and-stick approach would better leverage country $i$’s position in the trade system when dealing with its trade partner $j$. The trade-off is determining how much it should assist or contest $j$ without sacrificing its national interests and potentially increasing $\pi_{i}$.
@@
We can conceptualize $\pi_{i}$ as a container or pool within the dynamics of counterbalance equilibrium, where comparative and competitive advantages serve as the primary driving forces for the flows. In the inflow equation Eq. (4), country $i$ accumulates or absorbs authority from other countries through its direct influences. It derives more authority from influential trade partners (i.e., $\pi_{j}$ is large) than from non-influential ones, all else being equal. Additionally, it derives more authority from countries on which it has a significant direct influence (i.e., $P_{ji}$ is large), all else being equal. In Figure I, country $i$ imports goods from $j$ through the $i$th column of $P$, and thus, authority flows from $j$ to $i$ along the same column.
Conversely, country $i$ also contributes to other countries, as illustrated in the following outflow equation:
$\pi_{j}=\pi_{i}P_{ij}+\sum\limits_{k\neq i}\pi_{k}P_{kj}.$
The larger $P_{ij}$ or $\pi_{i}$, the more $i$ contributes to $j$’s authority $\pi_{j}$, all else being equal. In Figure I, country $i$ redistributes $\pi_{i}$ through the $i$th row of $P$, according to its shares of exports.
In the accounting of $\pi$, the two countries reach a break-even point when:
@@
Intuitively, if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, then country $i$ either takes more authority from or gives more to country $j$, respectively. For country $i$, a straightforward implication of Eq. (6) is to decrease or increase its imports from $j$ if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, respectively. It can also adjust its exports to $j$ accordingly. However, the outflow from $i$ to $j$ does not necessarily match the inflow. With the involvements of third parties, the inflows and outflows meet at the equilibrium Eq. (3) for $i$, i.e. $\pi_{i}P_{ij}+\sum\limits_{k\neq j}\pi_{i}P_{ik}=\pi_{j}P_{ji}+\sum\limits_{k\neq j}\pi_{k}P_{ki}=\pi_{i}$. For the same reason, it could be overly restrictive for all countries to balance their bilateral trades. For example, country $i$ can still maintain an overall zero net balance when it has a trade surplus with $j$ and a trade deficit with $k$ of the same amount.
Finally, Eq. (3) implies that $\pi$ is the unit-sum row eigenvector of the largest row eigenvalue of $P$. It represents a type of eigenvector centrality, a subclass of network centrality. In the literature, eigenvector centrality is the dominant eigenvector for an adjacency or relation matrix, which often has a zero diagonal (e.g., Bonacich, 1987). Our $P$ has a nonzero diagonal, allowing domestic and foreign goods to compete in the domestic market, and the eigenvector $\pi$ measures competitiveness (see Section 2.3). The same domestic products are used both domestically or overseas. Hu (2020) also retains a nonzero diagonal when any accepted student selects only one college to attend from multiple offers; thus, the selected college on the diagonal competes with the others that also accept the student and are off the diagonal. Furthermore, $P$ normalizes the scales of economies such that each row has a unit sum. Consequently, $P$ has the largest row eigenvalue $1$, common to all stochastic matrices.
2.3 $\pi$ as Competitive Advantages
The competitive advantage of country $i$ refers to its ability to outperform its competitors within the trade system. For any product exported from country $i$ to country $j$, the competitors of country $i$ include all other countries. Thus, its competitive advantage extends from bilateral country relations to a relationship with the entire system — it is the ability to surpass all other countries in the $j$th market for any $j\in\mathcal{N}$. In contrast, comparative advantage is an economy’s leverage of specialization, resource endowment, and technology over the trade partner’s opportunity cost; it is a bilateral relationship. Therefore, globalization may have made competitive advantages more prevalent than comparative advantages.
@@ 2.3 $\pi$ as Competitive Advantages
According to the literature (e.g., Porter, 1985; Stutz and Warf, 2010), competitive advantages rest on the direct control of competitive factors of production, such as access to rare resources or advanced technology, cheap but highly skilled labor, land, market affordability, and low cost of capital. These factors, imported from outside or manufactured domestically, are assembled in the final products. Thus, the imported goods carry the original producer’s competitiveness and should be attributed to the original maker. In the smartphone example, the competitive factors include the chips and patents from the USA, the labor from China, the precision manufacturing in Japan, and the sophisticated cameras from Korea. Various studies have also described qualitative strategies to enhance firm competitiveness. For example, Porter (1985) lists three generic strategies: cost leadership, differentiation, and focus.
@@
We find that $\pi$ quantifies national competitiveness. To introduce the quantitative competitiveness, we ignore the home bias in trade (cf., McCallum, 1995) and make a strong assumption for the time being: no bilateral trade deficit nor surplus. Denote $\tau_{j}$ as the $j$th row of $P$. Because of no bilateral surplus nor deficit, $\tau_{j}$ except for its $j$th element equals the imports from other countries to $j$, all in fractions of $j$’s production. So, it measures other countries’ aggregate comparative advantages over country $j$. Besides, all other countries also compete for the comparative advantages over $j$; a large value in $\tau_{j}$ means solid competitiveness to occupy the $j$th market and win a lion’s share of $j$’s imports. If we use the $j$th market as the reference to rank the competitiveness of all countries, then the competitiveness score is the vector $\tau_{j}$, which has a unit sum. Yet, the $j$th value is not excluded from $\tau_{j}$, because domestic products also compete with country $j$’s imports when consumers make their consumption choices, and they have no home bias.
We apply endogenous weighting to the references to combine these scores $\tau_{1},\tau_{2},\cdots,\tau_{n}$. These individual cardinal utilities are not comparable because it is more challenging to occupy a competitive market than a less competitive one. We make them comparable by weighting, and the weights are not exogenously assigned but endogenously determined. More weight should be placed on a more competitive reference market than on a less competitive one. Let the unknown weights be $\rho=(\rho_{1},\rho_{2},\cdots,\rho_{n})$ with $\sum\limits_{j=1}^{n}\rho_{j}=1$ and $\rho_{j}\geq 0$ for all $j\in\mathcal{N}$. Then, we have a weighted competitiveness score $\sum\limits_{j=1}^{n}\rho_{j}\tau_{j}=\rho P$. Both the weighted score and the weight vector quantify the competitiveness of the countries. To be consistent, the weighted average $\rho P$ should be a positive multiple of the weight vector $\rho$. By Theorem 1, $\rho$ is just $\pi$.
@@
In the endogenous weighting, worldwide consumers unconsciously participate in evaluating national competitiveness, encompassing all products and services. Essentially, countries as a whole unknowingly assess themselves without external interventions from business interests, mainstream media, advertising, governments, or even military threats. The structure of the global trade network is largely determined endogenously by the distribution of firm productivity within individual countries. Nevertheless, government policies can influence this distribution without significantly affecting the welfare of international trade by shifting imports or exports between countries. Policies can also enhance firm productivity through tax exemptions, subsidies, and imposing heavy tariffs on foreign products. However, national competitiveness $\pi_{i}$ is not merely the sum of the individual competitiveness of domestic firms; firm competitiveness are weighted in $\pi$ by consumers, wealthy or poor, domestic or overseas. Affordability thus plays a vital role in $\pi$. The positive weights imply a positive correlation between firm and national competitiveness.
@@
Additionally, $\pi$ represents a von Neumann-Morgenstern Bergson social welfare function, as a positively weighted sum of individual von Neumann-Morgenstern utility functions $\tau_{1},\tau_{2},\cdots,\tau_{n}$. Thus, the weighted utility function $\pi$ is determined not by the consumption of production factors or the added value, but by consumer preferences, judged independently of national interests or preferences. Pareto optimality is achieved when weighting these individual utilities: $\pi_{j}\geq\pi_{k}$ if $P_{ij}\geq P_{ik}$ for all $i\in\mathcal{N}$. The optimality remains invariant if any individual utility function is scaled by a positive constant or added by a constant, such as scaling $\tau_{i}$ by $g_{i}$. Harsanyi (1955) and Hammond (1987) provide additional ethical, risk, and utilitarianism connotations for this type of social welfare function. With the weights being $\pi$ itself, $\pi$ exhibits some special properties, discussed in Section 2.4.
@@
However, the social welfare function $\pi$ significantly departs from widely-used utility or welfare functions in the literature on trade wars or globalization (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014). It is not explicitly expressed in a specific functional form of production factors and total factor productivity but implied from the counterbalance equilibrium, as indirect imports also impact the final products or services. Exported goods and services have different utilities for exporting and importing countries. Therefore, each country has a unique utility function, and a single utility function cannot capture the diversity of all countries. We use fewer assumptions to accomplish $\pi$ by ignoring how efficiently each country packages production factors into final products and how differently it trades the products with others. It is precisely derived from the data and uses no parametric models, thus no unknown parameters need to be estimated, and no residuals exist to differentiate the data from a parametric model. Additionally, according to Barney and Felin (2013), micro-foundations research is still heavily debated, with management, strategy, and organization scholars having varying views on the link between macroeconomic phenomena and microeconomic agents’ behaviors. Due to broad firm and national heterogeneity, micro-founded strategic movements are particularly challenging on the network with flows of millions of products over tens of thousands of directional edges. With this complexity, a certain level of abstraction and ambiguity is necessary to grasp the essential parts of the network.
@@
The index $\pi_{i}$ contains a unique component, $\pi_{i}P_{ii}$, which does not originate from external source, as indicated by Eq. (4). This encompasses, but is not limited to, raw oil fields, natural gas reserves, rare minerals, and patented technologies, all of which remain integral to the nation’s core competitive factors. A global value chain typically involves finite steps but also includes indefinite domestic loops when tracing all the added values of the final products. For instance, in the smartphone example, the education, skills, and experience of manufacturing labor are ultimately developed domestically over many years, contributing to the final product. Additionally, numerous U.S. workers develop patents and design software domestically for extended periods. These inner loops affect other countries but not vice versa; this aspect of competitiveness never flows in from external sources. Consequently, social infrastructure, political institutions, monetary and fiscal policies, microeconomic environments, and other quality determinants that facilitate business all contribute to national competitiveness (e.g., Delgado et al., 2012).
@@
For a specific global value chain, at the other end of the unique component is the design and assembly of the final product. Product leadership bridges the market, consumers, strategic direction, technology innovation, sales, and production factors together; this leadership represents another type of competitiveness. Between the end product and the raw resources in the chain are a series of value-added processes for intermediate goods. Their competitiveness lies in cost efficiency or cost leadership, including the cost of financing. Technological advancements reduce labor costs, and public safety and social stability mitigate the risk of costs.
@@
Eq. (2) indicates a robust positive correlation between $\pi_{i}$ and $g_{i}$, the breadth of the economy. After adjusting for the sizes of $g_{i}$ in $P$, the interaction effects between competitiveness and economies of scale remain, thereby conferring some cost advantages. Nations export goods that are abundant and import goods that are scarce domestically. With significant abundance and shortage, country $i$ could have substantial nonzero elements in the $i$th column and row of the matrix $P$, resulting in considerable nonzero items summed in Eq. (2). A large economy also generates a high differentiation of goods and services, which offer location advantages and minimal transaction costs for domestic consumers. Free from tariffs and overseas shipping, these goods also possess a price advantage. Consequently, $P_{ii}$ is often the largest in the $i$th row of $P$, irrespective of home bias. Therefore, a vast economy with a comprehensively integrated industrial system (such as China and the USA) is likely, though not necessarily, to have a large value in $\pi$. A country with a single industrial structure is susceptible to intense competition, and its $\pi_{i}$ is fragile and volatile when competing with the USA for competitiveness.
@@
Hence, we use $\pi_{i}$ as an objective for country $i$ when formulating trade policies. It quantifies the economy’s depth and its interplay with breadth, encompassing both cooperation and non-cooperation with other nations. This measure is endogenously derived from the comparative advantages of all products and services. Given that each country has unique locations, natural and human resources, and economic systems, a single utility function that assumes one stereotype for all economies may be insufficient to capture this diversity. Our study acknowledges that each country, along with its trade stance with partners, is distinct. Each contributes inclusiveness value to the global economy, where international trade is a cornerstone.